OpenStudy (anonymous):

40% of what number is 20%

OpenStudy (amistre64):

If you mean 40% of what number is 20; the answer would be: part/% . 20/.4= 200/4 = 100/2 = 50. Double check: 20/50= 20(2)/50(2) = 40/100 = 40%.